# Structural and rigidity properties of Lie skew braces

### v1

@@ Abstract.
We investigate structural and rigidity properties of Lie skew braces (LSBs), objects essentially known in the literature as post–Lie groups, obtained by endowing a manifold with two compatible group laws that share the same identity element. LSBs extend skew left braces, which are central to the study of non-involutive set-theoretic solutions of the Yang–Baxter equation, to the smooth category. Our first main result (Theorem 1.1) shows that, for every connected LSB $(G,\cdot,\circ)$, linearity (in the simply-connected case) and solvability carry over from $(G,\cdot)$ to $(G,\circ)$, whereas the converse direction is rigid: if $(G,\circ)$ is nilpotent (respectively, semisimple) then $(G,\cdot)$ is forced to be solvable (respectively, isomorphic to $(G,\circ)$). Theorem 1.3 provides two “flexibility” statements: every non-linear simply connected Lie group $(G,\cdot)$ admits an LSB $(G,\cdot,\circ)$ such that $(G,\circ)$ is linear, and every simply connected solvable Lie group $(G,\circ)$ supports a LSB $(G,\cdot,\circ)$ such that $(G,\cdot)$ is nilpotent. A third result (Theorem 1.4) provides a complete existence table for non-trivial LSBs across the six standard Lie-group classes, abelian, nilpotent (non-abelian), solvable (non-nilpotent), simple, semisimple (non-simple) and mixed type, identifying precisely when a LSB can be built and when only the trivial or no structure occurs. Both the explicit constructions and the properties established in our theorems rely on a factorisation technique for Lie groups, on the correspondence between LSBs and regular subgroups of the affine group $\operatorname{Aff}(G,\cdot)$, which renders LSB theory equivalent to simply transitive affine actions, and on the theory of post–Lie algebras together with their integrability properties.
@@ 1. Introduction
where $a^{-1}$ denotes the inverse of $a$ in the group $(G,\cdot)$. LSBs are essentially just Post-Lie groups introduced in [2, Sections 3.3 and 4 ]. More precisely, the categories of LSBs and Post-Lie groups are equivalent (see Subsection 2.2 below). A LSB can be seen as a generalization of the skew left braces introduced by Guarnieri and Vendramin in [20] (see also [29]), originally developed to study non-involutive solutions of the Yang–Baxter equation [34]), and have since become a highly active area of research. Indeed, skew left braces correspond to LSB when $\dim(G)=0$ and $G$ is equipped with the discrete topology.
@@
    If $(G,\circ)$ is semisimple, then $(G,\cdot)\cong(G,\circ)$ (isomorphism of Lie groups).
@@
It is important to observe that the assumption of simple connectedness is essential for (S1) in Theorem 1.1 to hold (see Remark 4 below). Similarly, without the assumption of connectedness, (S2) fails. Indeed, there exist LSB structures $(G,\cdot,\circ)$ with $G$ disconnected such that $(G,\cdot)$ is solvable, but $(G,\circ)$ is not (see, for example, [28, Ex. 3.2]). In these examples, the group $G$ has countably many connected components. In the case where the LSB $(G,\cdot,\circ)$ has $G$ finite, i.e., $\dim G=0$ and $G$ consists of finitely many connected components, and hence we are in the realm of left skew braces, the question of whether the solvability of $(G,\cdot)$ implies the solvability of $(G,\circ)$ remains open. This conjecture was first posed by Leandro Vendramin and Agata Smoktunowicz in [29]. Partial progress on this problem has been made; see, for instance, [28], [14], and [31].
In [29] the authors show that if $(G,\cdot,\circ)$ is a finite left skew brace whose multiplicative group $(G,\cdot)$ is nilpotent, then the additive group $(G,\circ)$ is necessarily solvable. Combining this fact with property (S2) of Theorem 1.1 yields the following compact-group analogue (see Section 3 for a proof and Remark 5).
@@
    For any non-linear connected and simply-connected Lie group $(G,\cdot)$, there exists a LSB $(G,\cdot,\circ)$ such that $(G,\circ)$ is linear.
@@
    For any connected and simply-connected solvable Lie group $(G,\circ)$, there exists a LSB $(G,\cdot,\circ)$ such that $(G,\cdot)$ is nilpotent.
@@
The proofs of (R1) and (R2) of Theorem 1.1 are based on associating to a LSB $(G,\cdot,\circ)$ a post Lie algebra structure, namely a triple $(\mathfrak{g},[\cdot,\cdot]_{\cdot},[\cdot,\cdot]_{\circ})$, where $(\mathfrak{g},[\cdot,\cdot]_{\cdot})$ (resp. $(\mathfrak{g},[\cdot,\cdot]_{\circ})$) is the Lie algebra associated to $(G,\cdot)$ (resp. $(G,\circ)$) related by a post-Lie product (cf. Def. 2.7 below).
Not every post-Lie algebra arises as the infinitesimal counterpart of a LSB structure. Determining which PLAs can be integrated, i.e., obtained by differentiating a LSB, is a subtle and active area of research (see [2, 9, 11, 12, 18] and references therein). It is therefore of particular interest to construct non-trivial examples of LSBs.
In the following theorem, our third main result, we summarize the existence of non-trivial LSBs structures for various combinations of $(G,\cdot,\circ)$, as illustrated in the table. A check mark ($\checkmark$) indicates that a non-trivial LSB structure of the corresponding type exists, while a dash (--) that no such example is possible and a $\cong$ that the LSB is trivial, i.e. $(G,\cdot)\cong(G,\circ)$. Notice also, that to avoid overlaps between classes, we adopt the following conventions:
@@
To conclude, we stress how our results complement and extend, in the solvable and semisimple cases, the classification provided in [13, Th. 5.1], which concerns PLAs. This comparison allows for the immediate identification of several PLAs that cannot be integrated into a LSB. For example, there exist PLAs $(\mathfrak{g},[\cdot,\cdot]_{\cdot},[\cdot,\cdot]_{\circ})$ where $(\mathfrak{g},[\cdot,\cdot]_{\cdot})$ is a non-nilpotent solvable Lie algebra, while $(\mathfrak{g},[\cdot,\cdot]_{\circ})$ is reductive but non-abelian (see Example 5 below). According to condition (S2) in Theorem 1.1, such PLAs are not integrable into an LSB.
[…]
@@
The paper is organised as follows. Section 2 recalls the necessary background on Lie skew braces, post–Lie groups and algebras, and their realisation as regular subgroups of the affine group $\operatorname{Aff}(G,\cdot)$; in particular, it describes in detail the correspondence with simply transitive affine actions, introduces the $\lambda$–action, and sets up the integration functor that associates to each LSB its corresponding post–Lie algebra. Section 3 contains the proofs of Theorem 1.1-1.3 and Corollary 1.2: parts $\bf(S1)$-$\bf(S2)$ are obtained by lifting LSBs to the universal cover and exploiting the linearity of $\operatorname{Aff}_{0}(G,\cdot)$, while the rigidity statements $\bf(R1)$-$\bf(R2)$ follow from corresponding rigidity theorems for post-Lie algebras. The flexibility results and the existence table are derived through an explicit factorisation technique for Lie groups, the equivalence between LSBs and simply transitive affine actions (cf. Theorem 2.12 below), and integrability criteria for post-Lie algebrass
We would like to thank Giovanni Placini and Fabio Zuddas for carefully reading our paper and for their valuable comments and suggestions.
@@ 2.1. Lie skew braces
holds for all $a,b,c\in G$, where $a^{-1}$ denotes the inverse of $a$ in the group $(G,\cdot)$. From this it follows that the two groups also share they same identiity element denoted here by $e$. One can then define the category $\mathbf{LSB}$, whose objects are LSBs, and whose morphisms between two LSBs $(G_{1},\cdot_{1},\circ_{1})$ and $(G_{2},\cdot_{2},\circ_{2})$ are LSB homomorphisms, that is, smooth maps $\varphi:G_{1}\to G_{2}$ such that
@@
(cf. [20, Ex. 1.3 ]) Let $(G,\cdot)$ be a real Lie group. Define $a\circ b=a\cdot b$ for all $a,b\in G$. Then $(G,\cdot,\circ)$ is a LSBs. Similarly, defining $a\circ b=b\cdot a$ also yields a LSB structure on $G$. These LSBs are isomorphic if and only if $(G,\cdot)$ is abelian. For our aim these LSBs are not interesting since in both cases $(G,\cdot)\cong(G,\circ)$.
@@
is also a LSB, where the operations are defined componentwise by
@@
(cf. [20, Ex. 1.6]) Let $G$ be a real Lie group admitting a factorization $G=G_{1}G_{2}$, where $G_{1}$ and $G_{2}$ are Lie subgroups such that $G_{1}\cap G_{2}=\{1\}$. In particular, every element $a\in G$ can be uniquely written as $a=a_{1}\cdot a_{2}$ for some $a_{1}\in G_{1}$, $a_{2}\in G_{2}$. This factorization allows us to define a new group operation $\circ$ on the underlying manifold of $G$, as follows: for any $a,b\in G$, let $a=a_{1}\cdot a_{2}$, with $a_{1}\in G_{1}$ and $a_{2}\in G_{2}$ Then define
@@
With this operation, the triple $(G,\cdot,\circ)$ becomes a LSB. Indeed the map
@@
forms a LSB $(G_{1}\times G_{2},\cdot,\circ_{\alpha})$. Here, $(G_{1}\times G_{2},\circ_{\alpha})$ is precisely the semidirect product Lie group $G_{1}\rtimes_{\alpha}G_{2}$.
@@
where $\mathrm{Aut}(G,\cdot)$ acts on $G$ by evaluation; that is, for $f\in\mathrm{Aut}(G,\cdot)$ and $a\in G$, the action is given by evaluation.
@@
for all $a,b\in G$, $f,g\in\mathrm{Aut}(G,\cdot)$. Any subgroup $H$ of $\mathrm{Aff}(G,\cdot)$ acts smoothly on $G$:
@@
In particular $\mathrm{Aff}(G,\cdot)$ acts transitively on $G$ and the stabilizer of any $a\in G$ is a Lie subgroup of $\mathrm{Aff}(G,\cdot)$ isomorphic (as Lie group) to $G$.
@@
Let $(G,\cdot)$ be a connected Lie group. A subgroup $H\subseteq\mathrm{Aff}(G,\cdot)$ is said to be regular if for each $a\in G$ there exists a unique $(x,f)\in H$ such that $x\cdot f(a)=1$.
Notice that if $H$ is a regular subgroup of $\mathrm{Aff}(G,\cdot)$ then it is not hard to see that the projection $\pi_{1}:\mathrm{Aff}(G)\to G$ on the first factor, when restricted to $H$, is a bijection (see [20, Lemma 4.1]). It follows that $H$ is a closed subset of $\mathrm{Aff}(G,\cdot)$, and therefore $H$ is a Lie subgroup of $\mathrm{Aff}(G,\cdot)$. In fact, $H=({\pi_{1}}{|_{H}})^{-1}(G)$. Since $\pi_{1}$ is a continuous (even smooth) map and $G\subseteq\mathrm{Aff}(G,\cdot)\cong G\times\mathrm{Aut}(G,\cdot)$ is a closed submanifold, it follows that $H$ is closed as the preimage of a closed set under a continuous map.
Now, let us fix a LSB $(G,\cdot,\circ)$. Define the lambda-action
@@
Let $(G,\cdot,\circ)$ be a LSB. Then the map $\lambda$ is smooth.
@@
Now, since any Lie group is a locally compact topological space, it follows from [27, Theorem 46.11] that for each $a\in G$, the map
@@
Suppose $(G,\cdot,\circ)$ is a LSB. Define the regular Lie subgroup $H=\{(a,\lambda_{a})\mid a\in G\}\leq\mathrm{Aff}(G,\cdot)$, where $\lambda_{a}$ given by (3). The map
@@
where $(\pi_{1}|_{H})^{-1}(a)=(a,f)$. It follows that $(G,\circ)$ is a Lie group and (a direct computation shows that) $(G,\cdot,\circ)$ is a LSB.

### v3

@@ Abstract.
We investigate structural and rigidity properties of Lie skew braces (LSBs), objects essentially known in the literature as post–Lie groups, obtained by endowing a manifold with two compatible group laws that share the same identity element. LSBs extend skew left braces, which are central to the study of non-involutive set-theoretic solutions of the Yang–Baxter equation, to the smooth category. Our first main result (Theorem 1.1) shows that, for every connected LSB $(G,\cdot,\circ)$, linearity (in the simply-connected case) and solvability carry over from $(G,\cdot)$ to $(G,\circ)$, whereas the converse direction is rigid: if $(G,\circ)$ is nilpotent (respectively, semisimple) then $(G,\cdot)$ is forced to be solvable (respectively, isomorphic to $(G,\circ)$). Theorem 1.3 provides two “flexibility” statements: every non-linear simply connected Lie group $(G,\cdot)$ admits an LSB $(G,\cdot,\circ)$ such that $(G,\circ)$ is linear, and every simply connected solvable Lie group $(G,\circ)$ supports an LSB $(G,\cdot,\circ)$ such that $(G,\cdot)$ is nilpotent. A third result (Theorem 1.4) provides a complete existence table for non-trivial LSBs across the six standard Lie-group classes, abelian, nilpotent (non-abelian), solvable (non-nilpotent), simple, semisimple (non-simple) and mixed type, identifying precisely when an LSB can be built and when only the trivial or no structure occurs. Both the explicit constructions and the properties established in our theorems rely on a factorisation technique for Lie groups, on the correspondence between LSBs and regular subgroups of the affine group $\operatorname{Aff}(G,\cdot)$, which renders LSB theory equivalent to simply transitive affine actions, and on the theory of post–Lie algebras together with their integrability properties.
@@ 1. Introduction
where $a^{-1}$ denotes the inverse of $a$ in the group $(G,\cdot)$. LSBs are essentially just Post-Lie groups introduced in [2, Sections 3.3 and 4 ]. More precisely, the categories of LSBs and Post-Lie groups are equivalent (see Subsection 2.2 below). an LSB can be seen as a generalization of the skew left braces introduced by Guarnieri and Vendramin in [20] (see also [31]), originally developed to study non-involutive solutions of the Yang–Baxter equation [37]), and have since become a highly active area of research. Indeed, skew left braces correspond to LSB when $\dim(G)=0$ and $G$ is equipped with the discrete topology.
@@
    If $(G,\circ)$ is semisimple, then $(G,\cdot)$ and $(G,\circ)$ are locally isomorphic.
@@
It is important to observe that the assumption of simple connectedness is essential for (S1) in Theorem 1.1 to hold (see Remark 6 below). Similarly, without the assumption of connectedness, (S2) fails. Indeed, there exist LSB structures $(G,\cdot,\circ)$ with $G$ disconnected such that $(G,\cdot)$ is solvable, but $(G,\circ)$ is not (see, for example, [30, Ex. 3.2]). In these examples, the group $G$ has countably many connected components. In the case where the LSB $(G,\cdot,\circ)$ has $G$ finite, i.e., $\dim G=0$ and $G$ consists of finitely many connected components, and hence we are in the realm of left skew braces, the question of whether the solvability of $(G,\cdot)$ implies the solvability of $(G,\circ)$ remains open. This conjecture was first posed by Leandro Vendramin and Agata Smoktunowicz in [31]. Partial progress on this problem has been made; see, for instance, [30], [14], and [33].
In [31] the authors show that if $(G,\cdot,\circ)$ is a finite left skew brace whose additive group $(G,\cdot)$ is nilpotent, then the multiplicative group $(G,\circ)$ is necessarily solvable. Combining this fact with property (S2) of Theorem 1.1 yields the following compact-group analogue (see Section 3 for a proof and Remark 7).
@@
    For any non-linear connected and simply-connected Lie group $(G,\cdot)$, there exists an LSB $(G,\cdot,\circ)$ such that $(G,\circ)$ is linear.
@@
    For any connected and simply-connected solvable Lie group $(G,\circ)$, there exists an LSB $(G,\cdot,\circ)$ such that $(G,\cdot)$ is nilpotent.
@@
The proofs of (R1) and (R2) of Theorem 1.1 are based on associating to an LSB $(G,\cdot,\circ)$ a post Lie algebra structure, namely a triple $(\mathfrak{g},[\cdot,\cdot]_{\cdot},[\cdot,\cdot]_{\circ})$, where $(\mathfrak{g},[\cdot,\cdot]_{\cdot})$ (resp. $(\mathfrak{g},[\cdot,\cdot]_{\circ})$) is the Lie algebra associated to $(G,\cdot)$ (resp. $(G,\circ)$) related by a post-Lie product (cf. Def. 2.8 below).
Not every post-Lie algebra arises as the infinitesimal counterpart of an LSB structure. Determining which PLAs can be integrated, i.e., obtained by differentiating an LSB, is a subtle and active area of research (see [2, 9, 11, 12, 18] and references therein). It is therefore of particular interest to construct non-trivial examples of LSBs.
In the following theorem, our third main result, we summarize the existence of non-trivial LSBs structures for various combinations of $(G,\cdot,\circ)$, as illustrated in the table. A check mark ($\checkmark$) indicates that a non-trivial LSB structure of the corresponding type exists, while a dash (-) that no such example is possible and a $\cong$ that the LSB is trivial, i.e. $(G,\cdot)\cong(G,\circ)$. We write $\cong_{\mathrm{loc}}$ to denote local isomorphism of Lie groups, i.e. isomorphism of the corresponding Lie algebras. Notice also, that to avoid overlaps between classes, we adopt the following conventions:
@@
To conclude, we stress how our results complement and extend, in the solvable and semisimple cases, the classification provided in [13, Th. 5.1], which concerns PLAs. This comparison allows for the immediate identification of several PLAs that cannot be integrated into an LSB. For example, there exist PLAs $(\mathfrak{g},[\cdot,\cdot]_{\cdot},[\cdot,\cdot]_{\circ})$ where $(\mathfrak{g},[\cdot,\cdot]_{\cdot})$ is a non-nilpotent solvable Lie algebra, while $(\mathfrak{g},[\cdot,\cdot]_{\circ})$ is reductive but non-abelian (see Example 5 below). According to condition (S2) in Theorem 1.1, such PLAs are not integrable into an LSB.
Another interesting implication of our results concerns the following open question posed by the authors in [13], which relates to the existence of PLAs $(\mathfrak{g},[\cdot,\cdot]_{\cdot},[\cdot,\cdot]_{\circ})$ such that $(\mathfrak{g},[\cdot,\cdot]_{\cdot})$ is nilpotent and $(\mathfrak{g},[\cdot,\cdot]_{\circ})$ is perfect, i.e. $\mathfrak{g}=[\mathfrak{g},\mathfrak{g}]_{\circ}$. While we are currently not able to answer this question definitively, we can nevertheless state that any such PLA, if it exists, would not be integrable into an LSB due to condition (S2) in Theorem 1.1.
@@
The paper is organised as follows. Section 2 recalls the necessary background on Lie skew braces, post–Lie groups and algebras, and their realisation as regular subgroups of the affine group $\operatorname{Aff}(G,\cdot)$; in particular, it describes in detail the correspondence with simply transitive affine actions, introduces the $\lambda$-action, and sets up the derivation functor that associates to each LSB its corresponding post Lie algebra. Section 3 contains the proofs of Theorem 1.1-1.3 and Corollary 1.2: parts $\bf(S1)$-$\bf(S2)$ are obtained by lifting LSBs to the universal cover and exploiting the linearity of $\operatorname{Aff}_{0}(G,\cdot)$, while the rigidity statements $\bf(R1)$-$\bf(R2)$ follow from corresponding rigidity theorems for post-Lie algebras. The flexibility results and the existence table are derived through an explicit factorisation technique for Lie groups, the equivalence between LSBs and simply transitive affine actions (cf. Theorem 2.13 below), and integrability criteria for post-Lie algebras.
@@ 2.1. Lie skew braces
holds for all $a,b,c\in G$, where $a^{-1}$ denotes the inverse of $a$ in the group $(G,\cdot)$. From this it follows that the two groups also share the same identity element denoted here by $e$. One can then define the category $\mathbf{LSB}$, whose objects are LSBs, and whose morphisms between two LSBs $(G_{1},\cdot_{1},\circ_{1})$ and $(G_{2},\cdot_{2},\circ_{2})$ are LSB homomorphisms, that is, smooth maps $\varphi:G_{1}\to G_{2}$ such that
@@
(cf. [20, Ex. 1.3 ]) Let $(G,\cdot)$ be a real Lie group. Define $a\circ b=a\cdot b$ for all $a,b\in G$. Then $(G,\cdot,\circ)$ is an LSB. Similarly, defining $a\circ b=b\cdot a$ also yields an LSB structure on $G$. These LSBs are isomorphic if and only if $(G,\cdot)$ is abelian. For our aim these LSBs are not interesting since in both cases $(G,\cdot)\cong(G,\circ)$.
@@
is also an LSB, where the operations are defined componentwise by
@@
(cf. [20, Ex. 1.6]) Let $G$ be a real Lie group admitting a factorization $G=G_{1}G_{2}$, where $G_{1}$ and $G_{2}$ are Lie subgroups such that $G_{1}\cap G_{2}=\{1\}$. In particular, every element $a\in G$ can be uniquely written as $a=a_{1}\cdot a_{2}$ for some $a_{1}\in G_{1}$, $a_{2}\in G_{2}$. This factorization allows us to define a new group operation $\circ$ on the underlying manifold of $G$, as follows: for any $a,b\in G$, let $a=a_{1}\cdot a_{2}$, with $a_{1}\in G_{1}$ and $a_{2}\in G_{2}$ . Then define
@@
With this operation, the triple $(G,\cdot,\circ)$ becomes an LSB. Indeed the map
@@
forms an LSB $(G_{1}\times G_{2},\cdot,\circ_{\alpha})$. Here, $(G_{1}\times G_{2},\circ_{\alpha})$ is precisely the semidirect product Lie group $G_{1}\rtimes_{\alpha}G_{2}$.
@@
where $\mathrm{Aut}(G,\cdot)$ acts on $G$ by evaluation.
@@
for all $a,b\in G$, $f,g\in\mathrm{Aut}(G,\cdot)$. Any Lie subgroup $H$ of $\mathrm{Aff}(G,\cdot)$ acts smoothly on $G$:
@@
In particular $\mathrm{Aff}(G,\cdot)$ acts transitively on $G$ and the stabilizer of any $a\in G$ is a Lie subgroup of $\mathrm{Aff}(G,\cdot)$ isomorphic (as Lie group) to $\mathrm{Aut}(G,\cdot)$.
@@
Let $(G,\cdot)$ be a connected Lie group. A subgroup $H\subseteq\mathrm{Aff}(G,\cdot)$ is said to be regular if for each $a\in G$ there exists a unique $(x,f)\in H$ such that $x\cdot f(a)=e$, where $e$ denote the identity element of $(G,\cdot)$.
If $H$ is a regular subgroup of $\mathrm{Aff}(G,\cdot)$ then it is not hard to see that the projection $\pi_{1}:\mathrm{Aff}(G)\to G$ onto the first factor, when restricted to $H$, is a bijection (see [20, Lemma 4.1]).
Definition 2.3.
Let $(G,\cdot)$ be a connected Lie group. A subgroup $H\subseteq\mathrm{Aff}(G,\cdot)$ is said to be Lie regular if the following conditions are satisfied:
  1. (1)
    $H$ is a regular subgroup of $\mathrm{Aff}(G,\cdot)$;
  2. (2)
    $H$ is closed as a subspace of $\mathrm{Aff}(G,\cdot)$.
Remark 1.
By the Closed Subgroup Theorem, assumption (2) implies that $H$ is an embedded Lie subgroup of $\mathrm{Aff}(G,\cdot)$. In fact, one can weaken (2) to: “$H$ is an immersed Lie subgroup of $\mathrm{Aff}(G,\cdot)$’.’ Indeed, the action of $H$ on $G$,
$H\times G\longrightarrow G,\qquad(x,f)\cdot a:=x\,f(a),$
is smooth, free, and transitive. For a smooth transitive action of a Lie group $H$ on a Hausdorff smooth manifold $X$, the canonical map
$\Phi:\;H/H_{x}\longrightarrow X,\qquad\Phi(hH_{x})=h\cdot x,$
is a homeomorphism (in fact, a diffeomorphism once $H/H_{x}$ is given its standard smooth structure, see e.g. [26, Th. 21.18]). Applying this with $X=G$ and $x=e$, the stabilizer is $H_{e}=\{e\}$, hence
$\pi_{1}|_{H}:\;H\longrightarrow G$
is a homeomorphism, where $\pi_{1}:G\times\mathrm{Aut}(G,\cdot)\to G$ denotes the first projection under the identification $\mathrm{Aff}(G,\cdot)\cong G\rtimes\mathrm{Aut}(G,\cdot)$. Write $(\pi_{1}|_{H})^{-1}(a)=(a,\varphi_{a})$ and define
$\varphi:G\longrightarrow\mathrm{Aut}(G,\cdot),\qquad\varphi(a):=\varphi_{a}.$
Then $\varphi=\pi_{2}\circ(\pi_{1}|_{H})^{-1}$ is continuous (here $\pi_{2}$ is the second projection), and
$H=\{\,(a,\varphi(a)):a\in G\,\}$
is the graph of a continuous map into the Hausdorff group $\mathrm{Aut}(G,\cdot)$. Therefore $H$ is closed in $\mathrm{Aff}(G,\cdot)$.
Now, let us fix an LSB $(G,\cdot,\circ)$. Define the lambda-action
@@
Let $(G,\cdot,\circ)$ be an LSB. Then the map $\lambda$ is smooth.
@@
Now, since any Lie group is a locally compact topological space, it follows from [29, Theorem 46.11] that for each $a\in G$, the map
@@
Suppose $(G,\cdot,\circ)$ is an LSB. Define the regular Lie subgroup $H=\{(a,\lambda_{a})\mid a\in G\}\leq\mathrm{Aff}(G,\cdot)$, where $\lambda_{a}$ is given by (3). The map
@@
where $(\pi_{1}|_{H})^{-1}(a)=(a,f)$. It follows that $(G,\circ)$ is a Lie group and (a direct computation shows that) $(G,\cdot,\circ)$ is an LSB.